# Leveraging Spreading Activation for Improved Document Retrieval in Knowledge-Graph-Based RAG Systems

### v1

@@ 3.1 Benchmarks
We evaluate our system on two well-known MHQA benchmarks: MuSiQue [musique] and 2WikiMultiHopQA [2wiki].
One of the earliest and most widely used multi-hop QA datasets was HotpotQA [hp], created through crowdsourced multi-hop questions over Wikipedia articles. However, it has faced considerable criticism for not consistently requiring genuine multi-hop reasoning, as many questions could be answered through dataset artifacts or inference shortcuts [hp_criticism1, hp_criticism2, hp_criticsm3].
In response, 2WikiMultiHopQA [2wiki] refined the evaluation paradigm by combining unstructured Wikipedia text with structured Wikidata knowledge, thereby mitigating some of these issues. MuSiQue [musique] advanced this approach further by composing multi-hop questions in a bottom-up manner from a large pool of single-hop questions while enforcing strict connected reasoning. Therefore, in our experiments, we used 2WikiMultiHopQA and MuSiQue to evaluate the performance gains introduced by our methodology.
@@ 3.2 Baselines
In the experiments, we include three different types of baseline RAG methods against which we compare our proposed pipeline. First, we employ the Naive RAG baseline, which retrieves the top-$k$ documents with the highest cosine similarity between their embeddings and the query embedding, and feeds them directly into the LLM’s prompt together with the input query. We evaluate two variants of this baseline, with $k=5$ and $k=10$.
@@
Finally, we have a system based on query-decomposition methodology, where the LLM is prompted to decompose a complex question into a sequence of simpler, preferably single-hop queries, which are then used to sequentially gather context and derive intermediate answers.
@@ 4 Results
Two main observations can be made from the table. First, in experiments conducted with the phi4 model, SA-RAG alone outperformed all other baseline approaches on both benchmarks. In contrast, this result did not replicate with gemma3. Our inspection of the retrieved context in gemma3 experiments suggests that the absence of similar improvements was not due to missing or insufficient contextual information, but rather due to the model’s own reasoning limitations. Although phi4 is smaller, it was specifically designed for complex reasoning tasks, unlike gemma3.
The second observation is that SA-RAG provides consistent performance gains when introduced as a plug-and-play module in any of the examined training-free pipelines, in all of the experiment setups. As shown in the table, combining SA-RAG with CoT iterative retrieval yields the strongest performance, significantly outperforming all baseline RAG approaches, and providing 25% to 39% absolute gains in answer correctness compared to Naive RAG.
@@
Naive RAG 0.25 0.36 45 0.26 0.36 44
Naive RAG (k=10) 0.23 0.38 48 0.32 0.41 53
CoT RAG 0.33 0.44 55 0.39 0.50 58
CoT RAG (k=10) 0.33 0.46 55 0.34 0.45 56
Query-decomposition 0.34 0.48 55 0.40 0.51 60
@@ 6 Conclusion & Limitation
In this paper, we addressed the challenges of document retrieval in RAG systems for complex tasks that require multi-step reasoning and evidence aggregation from multiple documents, such as MHQA. We proposed integrating the SA information retrieval algorithm with RAG approaches based on knowledge graphs to overcome the limitations of current systems without the need for expensive language model fine-tuning. Our experiments demonstrate the potential of this technique by achieving, with a single retrieval step, performance equivalent to or better than several training-free RAG baselines that use iterative retrieval, and further show that it can be combined with methods like CoT to significantly boost system performance, even when using small open-weight language models.

### v2

@@ 3.1 Benchmarks
We evaluate our system on two well-known MHQA benchmarks: MuSiQue [40] and 2WikiMultiHopQA [20].
One of the earliest and most widely used multi-hop QA datasets was HotpotQA [47], created through crowdsourced multi-hop questions over Wikipedia articles. However, it has faced considerable criticism for not consistently requiring genuine multi-hop reasoning, as many questions could be answered through dataset artifacts or inference shortcuts [5, 38, 29].
In response, 2WikiMultiHopQA [20] refined the evaluation paradigm by combining unstructured Wikipedia text with structured Wikidata knowledge, thereby mitigating some of these issues. MuSiQue [40] advanced this approach further by composing multi-hop questions in a bottom-up manner from a large pool of single-hop questions while enforcing strict connected reasoning. Therefore, in our experiments, we used 2WikiMultiHopQA and MuSiQue to evaluate the performance gains introduced by our methodology.
@@ 3.2 Baselines
In the experiments, we include four different types of baseline RAG methods against which we compare our proposed pipeline.
First, we employ the Naive RAG baseline, which retrieves the top-$k$ documents with the highest cosine similarity between their embeddings and the query embedding, and feeds them directly into the LLM’s prompt together with the input query. We evaluate two variants of this baseline, with $k=5$ and $k=10$.
@@
Third, we have a system based on query-decomposition methodology, where the LLM is prompted to decompose a complex question into a sequence of simpler, preferably single-hop queries, which are then used to sequentially gather context and derive intermediate answers.
Finally, we use HippoRAG as a graph-based baseline, specifically the second version proposed in [16], for which state-of-the-art performance is reported on the MHQA benchmarks used in our evaluation.
@@ 4 Results
Naive RAG 0.25 0.36 46 0.26 0.36 45
Naive RAG (k=10) 0.23 0.38 49 0.32 0.41 54
CoT RAG 0.33 0.44 56 0.39 0.50 58
CoT RAG (k=10) 0.33 0.46 56 0.34 0.45 56
Query-decomposition 0.34 0.48 56 0.40 0.51 61
HippoRAG 2 0.32 0.52 70 0.31 0.47 59
@@
Two main observations emerge from the table. First, in experiments with the phi4 model, SA-RAG alone outperformed all iterative baselines on both benchmarks and achieved performance comparable to HippoRAG 2, with slightly lower correctness on MuSiQue. With gemma3, however, HippoRAG 2 outperformed SA-RAG on both benchmarks, while the query-decomposition and CoT baselines outperformed it on MuSiQue. Our inspection of the retrieved context in the gemma3 experiments suggests that the absence of similar improvements was not due to missing or insufficient contextual information, but rather to limitations in the model’s reasoning. Although phi4 is smaller, it was specifically designed for complex reasoning tasks, unlike gemma3.
The second observation is that SA-RAG provides consistent performance gains when introduced as a plug-and-play module in any of the examined iterative pipelines, in all of the experiment setups. As shown in the table, combining SA-RAG with CoT retrieval yields the strongest performance, substantially outperforming all baseline RAG approaches, providing up to 10% absolute gain in answer correctness over HippoRAG 2, and 25% to 39% relative to Naive RAG.
@@ 6 Conclusion & Limitation
In this paper, we addressed the challenges of document retrieval in RAG systems for complex tasks that require multi-step reasoning and evidence aggregation from multiple documents, such as MHQA. We proposed integrating the SA information retrieval algorithm with RAG approaches based on heterogeneous knowledge graphs to overcome the limitations of current systems without the need for expensive language model fine-tuning. Our experiments demonstrate the potential of this technique by achieving, with a single retrieval step, performance equivalent to or better than several advanced training-free baselines, and further show that it can be combined with methods like CoT to significantly boost system performance, even when using small open-weight language models.